# Envelopes in the class of Banach algebras of polynomial growth and $C^{\infty}$-functions of a finite number of free variables

### v4

@@ 4. Envelopes of certain non-commutative algebras
It follows that the upper right entries of $\widetilde{\pi}_{p,\lambda}(u)^{r}$ and $\widetilde{\pi}^{\prime}_{p,\mu}(u)^{r}$ vanish when $r\neq p$. Putting $r=p-1$, we can write $\rho^{\prime}$ as the product of maps $\rho^{\prime}_{r}\!:C^{\infty}(\Omega)\to C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})$. The explicit form of $\rho^{\prime}_{r}$ is as follows. Let elements be written as in (4.1) and note that $\widetilde{\pi}_{r+1,\lambda}(f_{r})$ is the diagonal matrix with the entries $f_{r}(\lambda q^{r}),\ldots,f_{r}(\lambda)$ and $\widetilde{\pi}_{r+1,\lambda}(g_{r})$ has $g_{r}(0)$ at each diagonal entry. Applying the equality $f_{r}(0)=g_{r}(0)$, we conclude that for every $r$,
$\rho^{\prime}_{r}(f_{r},g_{r})=(\lambda\mapsto C_{q}\lambda^{q}f_{r}(\lambda q^{r})+f_{r}(0),\,\mu\mapsto C^{\prime}_{q}\mu^{q}g_{r}(\mu q^{r})+g_{r}(0)),$
where $C_{q}$ and $C^{\prime}_{q}$ are non-zero numbers depending only on $q$.
So $\rho^{\prime}_{r}$ is the restriction to $C^{\infty}(\Omega)$ of the product of two continuous linear endomorphism of $C^{\infty}(\mathbb{R})$ each of which is a composition of three linear maps, the first is induced by a diffeomorphism, the second is the multiplication operator on a polynomial and the third is a shift by a one-dimension operator. Each of these maps has closed range, the second map by Proposition 3.6 and the first and third because they are topological isomorphisms. Also, it is easy to see that all three maps are injective. Thus, by the inverse mapping theorem for Fréchet spaces, they are topologically injective and so is $\rho^{\prime}$ by Part (A) of Lemma 4.4. This completes the proof of Part (A).
@@
Note that $\varphi(u)$ is nilpotent by Lemma 4.2. Then $\mathbb{R}[u]\to B\!:u\mapsto\varphi(u)$ can be extended to continuous linear map $\mathbb{R}[[u]]\to B$ (this a partial case of Theorem 4.5). Take the composition of the tensor product of these two maps and the linearization of the multiplication in $B$,
@@
(B) It is easy to check by induction that $(ba)^{n}a^{-n}=q^{n(n-1)/2}b^{n}$ for every $n\in\mathbb{N}$. By Lemma 4.2, $ba$ is nilpotent and therefore is $b$. Similarly, we obtain that $c$ is also nilpotent. ∎
@@
Denote by $A_{q}$ the universal real associative algebra with generators $a$, $a^{-1}$, $b$, $c$ and the relations in (4.5). Combining Lemmas 4.6 and 4.7 we get the following assertion.
@@
The relations in (4.5) easily implies that
@@
Taking the composition of the inverse to this map and the tensor product of natural embeddings $\mathbb{R}[a,a^{-1}]\to C^{\infty}(\mathbb{R}^{\times})$ and $\mathbb{R}[b,c]\to\mathbb{R}[[b,c]]$ we obtain an injective linear map
@@
We can identify $C^{\infty}(\mathbb{R}^{\times})\mathbin{\widehat{\otimes}}\mathbb{R}[[b,c]]$ with a power of $C^{\infty}(\mathbb{R}^{\times})$, where the copies labelled by $\mathbb{Z}_{+}^{2}$. Then $\widetilde{\pi}_{p}$ maps a double sequence $(f_{ij})$ to the matrix in (4.9). Partition $\mathbb{Z}_{+}^{2}$ into the union of subsets given by the condition $i+j=r$ and write $C^{\infty}(\mathbb{R}^{\times})\mathbin{\widehat{\otimes}}\mathbb{R}[[b,c]]$ as $\prod_{r=0}^{\infty}C^{\infty}(\mathbb{R}^{\times})^{r}$. Consider the map
@@
We begin with formulating a result from [5] that was the starting point for writing this article. Recall that a finite-dimensional real Lie algebra $\mathfrak{g}$ is said to be triangular if it is solvable and for every $x\in\mathfrak{g}$ all the eigenvalues of the linear operator $\mathop{\mathrm{ad}}\nolimits x$ belong to $\mathbb{R}$.
@@
Consider the corresponding PBW basis $\{e^{\alpha}\!:=e_{1}^{\alpha_{1}}\cdots e_{m}^{\alpha_{m}}\!:\,\alpha\in\mathbb{Z}_{+}^{m}\}$ in $U(\mathfrak{g})$ and the linear map $U(\mathfrak{g})\to C^{\infty}_{\mathfrak{g}}$ with dense image given by identifying $e_{1},\ldots,e_{k}$ with the coordinate functions on $\mathbb{R}^{k}$ and the embedding $\mathbb{R}[e_{k+1},\ldots,e_{m}]\to\mathbb{R}[[e_{k+1},\ldots,e_{m}]]$. The following result combines Theorems 4.3 and 4.4 in [5].
@@
The main idea of the proof is the same as in Theorems 4.3 and 4.10 but the reasoning is cumbersome and contains many technicalities; for details see [5].
@@
It is the simplest example of a solvable, but not triangular, Lie algebra (the operator $\mathop{\mathrm{ad}}\nolimits x_{1}$ has eigenvalues $i,0,-i$). It is easy to see that $[\mathfrak{e}_{2},\mathfrak{e}_{2}]$ is a minimal ideal with basis $x_{2},x_{3}$. Thus $\mathfrak{e}_{2}^{\mathbf{tri}}\cong\mathfrak{e}_{2}/[\mathfrak{e}_{2},\mathfrak{e}_{2}]$. Then $\widehat{U}(\mathfrak{e}_{2})^{\,\mathsf{PG}}$ is isomorphic to $C^{\infty}(\mathbb{R})$ by Theorem 4.14.
@@ 5. Envelopes of free algebras
We need another version of tensor algebra — formal tensor algebra.
@@
Let $E$ be a complete (real or complex) locally convex space. The formal tensor algebra of $E$ is a real Arens–Michael algebra $[T](E)$ with a continuous linear map $\mu\!:E\to[T](E)$ satisfying the following condition. For every Banach algebra $B$ and every continuous linear map $\psi\!:E\to B$ with nilpotent range there is a unique unital continuous homomorphism $\widehat{\psi}\!:[T](E)\to B$ such that the diagram
@@
Let $E$ be a complete (real or complex) locally convex space. The direct product of all tensor products of the form $E\mathbin{\widehat{\otimes}}\cdots\mathbin{\widehat{\otimes}}E$, including $\mathbb{R}$ (or $\mathbb{C}$), with multiplication extended from the usual tensor algebra $T(E)$ is the formal tensor algebra of $E$.
@@
It is easy to see that $\Pi\!:=\prod_{k=0}^{\infty}E^{\mathbin{\widehat{\otimes}}k}$ is a unital Arens–Michael algebra with respect to the multiplication induced by the isomorphisms $E^{\mathbin{\widehat{\otimes}}k}\mathbin{\widehat{\otimes}}E^{\mathbin{\widehat{\otimes}}l}\cong E^{\mathbin{\widehat{\otimes}}(k+l)}$, $k,l\in\mathbb{Z}_{+}$.
@@
Let $k\geqslant 2$ and $e_{1},\ldots,e_{k}$ denote the generators of $\mathfrak{f}_{k}$. Then $[\mathfrak{f}_{k},\mathfrak{f}_{k}]$ is an infinitely generated free Lie algebra with the generators
$[e_{j_{1}},[e_{j_{2}},[\cdots[e_{j_{s-1}},e_{j_{s}}]\cdots]],\qquad j_{1}\geqslant j_{2}\geqslant\cdots j_{s-1}<j_{s},\,s\geqslant 2.$ (5.1)
@@
Denote by $\Delta$ the set of all pairs $(l,j)$ of positive integers such that $k\geqslant l>j\geqslant 1$. Here and throughout this section, when using the notation $l$ and $j$, it is always assumed that they form a pair $\delta\in\Delta$ and vice versa. It is natural to use polynomials in the operators $\mathop{\mathrm{ad}}\nolimits e_{i}$. For $\delta\in\Delta$ and $\beta=(\beta_{1},\ldots,\beta_{l})\in\mathbb{Z}_{+}^{l}$ put
$g_{\delta,\beta}\!:=(\mathop{\mathrm{ad}}\nolimits{e_{1}})^{\beta_{1}}\cdots(\mathop{\mathrm{ad}}\nolimits{e_{l-1}})^{\beta_{l-1}}(\mathop{\mathrm{ad}}\nolimits{e_{l}})^{\beta_{l}+1}(e_{j}).$ (5.2)
@@
where $T(V)$ is a tensor algebra. Note that (5.3) also holds for $k=1$ since in this case $\Delta$ is empty and so $V=0$.
@@
$\Psi_{\delta}\!:\mu_{1}^{\beta_{1}}\cdots\mu_{l}^{\beta_{l}}\mapsto\mathop{\mathrm{ad}}\nolimits e_{1}^{\beta_{1}}\cdots\mathop{\mathrm{ad}}\nolimits e_{l}^{\beta_{l}}$ (5.4)
@@
for $k\in\mathbb{N}$. (Here $[T](C^{\infty}(X))$ is the formal tensor algebra associated with the Fréchet space $C^{\infty}(X)$; see the beginning of the section.)
@@
If an element $b$ of a real Banach algebra $B$ is of polynomial growth, then so is the operator $\mathop{\mathrm{ad}}\nolimits b\!:B\to B$.
@@
where the hat means the Fourier transform. By Lemma 5.5, the operators $\mathop{\mathrm{ad}}\nolimits b_{1},\ldots,\mathop{\mathrm{ad}}\nolimits b_{k}$ are of polynomial growth. Therefore the function
$(s_{1},\ldots,s_{l})\mapsto\|\widehat{f}(s_{1},\ldots,s_{l})\,\exp\bigl{(}is_{1}\mathop{\mathrm{ad}}\nolimits b_{1}\bigr{)}\cdots\exp\bigl{(}is_{l}\mathop{\mathrm{ad}}\nolimits b_{l}\bigr{)}([b_{l},b_{j}])\|$
is absolutely integrable. Thus, the integral exists not only in the weak sense (of Gelfand and Pettis), but also in the strong sense (of Bochner) [22, p. 80, Theorem 3.7.4]. Applying Theorem 4.5 again, now to the tuple $(\mathop{\mathrm{ad}}\nolimits b_{1},\ldots,\mathop{\mathrm{ad}}\nolimits b_{k})$, we see that $\Psi_{\delta}$ is indeed an extension of the ordered calculus in (5.4). From the definition of $C^{\infty}(X)$ (see (5.5)), we have an extension of the map $V\to B$ to $C^{\infty}(X)\to B$.
@@
$\exp\bigl{(}is_{l}\mathop{\mathrm{ad}}\nolimits b_{l}\bigr{)}([b_{l},b_{j}])=[\exp\bigl{(}is_{l}b_{l})\,b_{l}\exp\bigl{(}-is_{l}b_{l})\,,\,\exp\bigl{(}is_{l}b_{l})\,b_{j}\exp\bigl{(}-is_{l}b_{l})].$
By Theorem 2.4, each Banach algebra of polynomial growth is commutative modulo the radical. Therefore the integrand in (5.8) belongs to $\mathop{\mathrm{Rad}}\nolimits B$ for all $s_{1},\ldots,s_{l}$. Since the radical of a Banach algebra is closed and the integral converges in the strong sense, we have that $\Psi_{\delta}(f)\in\mathop{\mathrm{Rad}}\nolimits B$ for every $f\in C^{\infty}(X_{\delta})$. Therefore the image of $C^{\infty}(X)$ is also in $\mathop{\mathrm{Rad}}\nolimits B$.
By Theorem 2.4, $\mathop{\mathrm{Rad}}\nolimits B$ is nilpotent. Then we can apply the universal property for $[T](C^{\infty}(X))$ in Definition 5.1 and extend the continuous linear map $C^{\infty}(X)\to B$ constructed above to a continuous linear map from $[T](C^{\infty}(X))$ to $B$. The claim is proved.
@@
The main idea of the proof of is the same as for Theorem 4.3 in [5] and Theorems 4.3 and 4.10 in this paper. We endow $U(\mathfrak{f}_{k})$ with the topology inherited from $C^{\infty}_{\mathfrak{f}_{k}}$ and construct a topologically injective homomorphism from $U(\mathfrak{f}_{k})$ to a product of algebras of the form $C^{\infty}(M,\mathrm{T}_{p})$, where $M$ is a manifold.
@@
$g_{\beta}=(\mathop{\mathrm{ad}}\nolimits{e_{1}})^{\beta_{1}}(\mathop{\mathrm{ad}}\nolimits{e_{2}})^{\beta_{2}+1}(e_{1}).$ (5.14)
@@
Then, applying inductively (5.13) with (5.12) and next (5.11) and using the fact that $\mathfrak{c}_{2}$ is a Lie ideal, we get the result. ∎
@@
In the proof of Theorem 5.6 we also need another lemma on topologically injective maps. First note that a family $(\psi_{ij}\!:X_{i}\to Y_{j};\,i\in I,\,j\in J)$ of continuous linear maps of complete locally convex spaces induces a continuous linear map $\prod_{i\in I}X_{i}\to\prod_{j\in J}Y_{j}$ under the assumption that for every $j$ the subset of $i$ such that $\psi_{ij}\neq 0$ is finite. Indeed, the finiteness condition implies that there is a morphism from $\prod_{i\in I}X_{i}$ to $Y_{j}$ for every $j$ and hence from $\prod_{i\in I}X_{i}$ to $\prod_{j\in J}Y_{j}$.
Lemma 5.12.
Let $(\psi_{nm}\!:X_{n}\to Y_{m};\,n,m\in\mathbb{N})$ be a family of continuous linear maps of Fréchet spaces such that $\psi_{nm}=0$ for $n>m$ and $\psi_{mm}$ is topologically injective for every $m$. Then the map $\prod X_{n}\to\prod Y_{m}$ induced by this family is topologically injective.
Proof.
It follows from a theorem of Dugundji [18, Theorem 4.1] that a continuous linear map between Fréchet spaces is topologically injective if and only if it has a continuous left inverse map (possibly, non-linear).
is left inverse (cf. the formula for the inverse of a triangle $2\times 2$ matrix) and so we have the topological injectivity.
Further, it easily follows by induction that $\prod_{n\leqslant N}X_{n}\to\prod_{m\leqslant N}Y_{m}$ also has a continuous left inverse map for every $N\in\mathbb{N}$. Moreover, the sequence of left inverse maps is compatible with projections $\prod^{N+1}X_{n}\to\prod^{N}X_{n}$ and $\prod^{N+1}Y_{m}\to\prod^{N}Y_{m}$. Since an infinite product is a projective limit of finite products (see, e.g., [1, Exercise 11B]), we have a continuous map $\prod Y_{m}\to\prod X_{n}$ that is, as easy to see, left inverse to $\prod X_{n}\to\prod Y_{m}$. Thus the latter map is topologically injective. ∎

### v5

@@ 3. Envelopes of certain commutative and non-commutative algebras
Let an element $a$ of $\mathcal{R}(\mathbb{R}^{2}_{q})$ be written as in (3.1). To find the upper right entries of $\widetilde{\pi}_{p,\lambda}(a)$ and $\widetilde{\pi}^{\prime}_{p,\mu}(a)$ note that $\widetilde{\pi}_{p,\lambda}(u)^{k}$ and $\widetilde{\pi}^{\prime}_{p,\mu}(u)^{k}$ vanish when $k\geqslant p$. Considering the case where $r\leqslant p-1$, we see that the only one non-trivial entry in the right column of $\widetilde{\pi}_{p,\lambda}(u)^{k}$ and $\widetilde{\pi}^{\prime}_{p,\mu}(u)^{k}$ equals respectively $\lambda^{k}q^{k(k+1)/2}$ and $\mu^{k}q^{-k(k-1)/2}$ both at the $(p-1-k)$th place.
Note that $\widetilde{\pi}_{p,\lambda}(\Phi_{x}(f_{j}))$ is the diagonal matrix with the entries $f_{j}(\lambda q^{p-1}),\ldots,f_{j}(\lambda)$ and
$\widetilde{\pi}_{p,\lambda}(\Phi_{y}(g_{j}))=\begin{pmatrix}g_{j}(0)&g_{j}^{\prime}(0)&0&\ldots&g_{j}^{(p-1)}(0)/(p-1)!\\ 0&g_{j}(0)&g_{j}^{\prime}(0)&\ldots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&0&\ddots&\ddots&g_{j}^{\prime}(0)\\ 0&0&\ldots&0&g_{j}(0)\end{pmatrix}.$
Therefore the upper right entry of $\widetilde{\pi}_{p,\lambda}(a)$ is equal to
$f_{p-1}(\lambda q^{p-1})\lambda^{p-1}q^{p(p-1)/2}+\sum_{k=0}^{p-1}\frac{g^{(p-1-k)}_{k}(0)}{(p-1-k)!}\lambda^{k}q^{k(k+1)/2}.$ (3.5)
Similarly, we have that the upper right entry of $\widetilde{\pi}^{\prime}_{p,\mu}(a)$ is equal to
$g_{p-1}(\mu q^{1-p})\mu^{p-1}q^{-(p-1)(p-2)/2}+\sum_{k=0}^{p-1}\frac{f^{(p-1-k)}_{k}(0)}{(p-1-k)!}\mu^{k}q^{-k(k-1)/2}.$ (3.6)
By continuity, (3.5) and (3.6) hold for every element of $C^{\infty}(\mathbb{R}^{2}_{q})$. We can treat $\rho^{\prime}$ as the map induced by a family $(\psi_{p-1,k}:X_{p-1}\to Y_{k};p,k\in\mathbb{N})$ of continuous linear maps, where $X_{p-1}=C^{\infty}(\Omega)$ and $Y_{k}=C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})$. Using the fact that $f_{p-1}(0)=g_{p-1}(0)$, we conclude that $\psi_{p-1,p}$ maps $((f_{p-1},g_{p-1}))$ to
$(\lambda\mapsto(f_{p-1}(\lambda q^{p-1})+f_{p-1}(0))\lambda^{p-1}q^{p(p-1)/2},\,\mu\mapsto(g_{p-1}(\mu q^{1-p})+g_{p-1}(0))\mu^{p-1}q^{-(p-1)(p-2)/2}).$
Also, $\psi_{p-1,k}=0$ when $k>p-1$. Lemma 3.11 implies that to prove that $\rho^{\prime}$ is topologically injective it suffices to prove that each of $\psi_{p-1,p}$ is topologically injective. We can assume that $\psi_{p-1,p}$ is defined on the whole $C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})$. Thus it is sufficient to check the topological injectivity of two maps on $C^{\infty}(\mathbb{R})$.
Each of them is a composition of three linear maps, the first is induced by a diffeomorphism, the second is a shift by a one-dimension operator and the third is the multiplication operator on a polynomial. Each of these maps has closed range — the first and second maps because they are topological isomorphisms and the third map by Proposition 3.9. Also, it is easy to see that all three maps are injective. Thus, by the inverse mapping theorem for Fréchet spaces, they are topologically injective and thus this completes the proof of Part (A).
@@
Note that $\varphi(u)$ is nilpotent by Lemma 3.7. Then $\mathbb{R}[u]\to B\!:u\mapsto\varphi(u)$ can be extended to continuous linear map $\mathbb{R}[[u]]\to B$ (this is a partial case of Theorem 3.12). Take the composition of the tensor product of these two maps and the linearization of the multiplication in $B$,
@@
(B) It is easy to check by induction that $(ba)^{n}a^{-n}=q^{n(n-1)/2}b^{n}$ for every $n\in\mathbb{N}$. By Lemma 3.7, $ba$ is nilpotent and hence so is $b$. Similarly, we obtain that $c$ is also nilpotent. ∎
@@
Denote by $A_{q}$ the universal real associative algebra with generators $a$, $a^{-1}$, $b$, $c$ and the relations in (3.7). Combining Lemmas 3.13 and 3.14 we get the following assertion.
@@
The relations in (3.7) easily imply that
@@
Taking the composition of the inverse to the previous map and the tensor product of natural the embeddings $\mathbb{R}[a,a^{-1}]\to C^{\infty}(\mathbb{R}^{\times})$ and $\mathbb{R}[b,c]\to\mathbb{R}[[b,c]]$ we obtain an injective linear map
@@
We can identify $C^{\infty}(\mathbb{R}^{\times})\mathbin{\widehat{\otimes}}\mathbb{R}[[b,c]]$ with a power of $C^{\infty}(\mathbb{R}^{\times})$, where the copies labelled by $\mathbb{Z}_{+}^{2}$. Then $\widetilde{\pi}_{p}$ maps a double sequence $(f_{ij})$ to the matrix as in (3.11). Partition $\mathbb{Z}_{+}^{2}$ into the union of subsets given by the condition $i+j=r$ and write $C^{\infty}(\mathbb{R}^{\times})\mathbin{\widehat{\otimes}}\mathbb{R}[[b,c]]$ as $\prod_{r=0}^{\infty}C^{\infty}(\mathbb{R}^{\times})^{r}$. Consider the map
@@
We begin with formulating a result from [5] that was the starting point for writing this article. Recall that a finite-dimensional real Lie algebra $\mathfrak{g}$ is said to be triangular if it is solvable and for every $x\in\mathfrak{g}$ all the eigenvalues of the linear operator $\operatorname{ad}x$ belong to $\mathbb{R}$.
@@
Consider the corresponding Poincaré–Birkoff–Witt basis $\{e^{\alpha}\!:=e_{1}^{\alpha_{1}}\cdots e_{m}^{\alpha_{m}}\!:\,\alpha\in\mathbb{Z}_{+}^{m}\}$ in $U(\mathfrak{g})$ and the linear map $U(\mathfrak{g})\to C^{\infty}_{\mathfrak{g}}$ with dense image given by identifying $e_{1},\ldots,e_{k}$ with the coordinate functions on $\mathbb{R}^{k}$ and the embedding $\mathbb{R}[e_{k+1},\ldots,e_{m}]\to\mathbb{R}[[e_{k+1},\ldots,e_{m}]]$. The following result combines Theorems 4.3 and 4.4 in [5].
@@
The main idea of the proof is the same as in Theorems 3.8 and 3.17 but the reasoning is cumbersome and contains many technicalities; for details see [5].
@@
It is the simplest example of a solvable, but not triangular, Lie algebra (the operator $\operatorname{ad}x_{1}$ has eigenvalues $i,0,-i$). It is easy to see that $[\mathfrak{e}_{2},\mathfrak{e}_{2}]$ is a minimal ideal with basis $x_{2},x_{3}$. Thus $\mathfrak{e}_{2}^{\mathbf{tri}}\cong\mathfrak{e}_{2}/[\mathfrak{e}_{2},\mathfrak{e}_{2}]$. Then $\widehat{U}(\mathfrak{e}_{2})^{\,\mathsf{PG}}$ is isomorphic to $C^{\infty}(\mathbb{R})$ by Theorem 3.21.
@@ 4. Envelopes of free algebras
We need another version of tensor algebra — formal tensor algebra; cf. [16] in the Banach space case.
@@
Let $E$ be a complete (real or complex) locally convex space. The formal tensor algebra of $E$ is an Arens–Michael algebra $[T](E)$ with a continuous linear map $\mu\!:E\to[T](E)$ satisfying the following condition. For every Banach algebra $B$ and every continuous linear map $\psi\!:E\to B$ with nilpotent range there is a unique unital continuous homomorphism $\widehat{\psi}\!:[T](E)\to B$ such that the diagram
@@
Let $E$ be a complete (real or complex) locally convex space. The direct product of all projective tensor products of the form $E\mathbin{\widehat{\otimes}}\cdots\mathbin{\widehat{\otimes}}E$, including $\mathbb{R}$ (or $\mathbb{C}$), with multiplication extended from the usual tensor algebra $T(E)$ is the formal tensor algebra of $E$.
@@
It is easy to see that $\Pi\!:=\prod_{k=0}^{\infty}E^{\mathbin{\widehat{\otimes}}k}$ is a unital Arens–Michael algebra with respect to the multiplication induced by the isomorphisms $E^{\mathbin{\widehat{\otimes}}k}\mathbin{\widehat{\otimes}}E^{\mathbin{\widehat{\otimes}}l}\cong E^{\mathbin{\widehat{\otimes}}(k+l)}$; $k,l\in\mathbb{Z}_{+}$.
@@
Let $k\geqslant 2$ and let $e_{1},\ldots,e_{k}$ denote the generators of $\mathfrak{f}_{k}$. Then $[\mathfrak{f}_{k},\mathfrak{f}_{k}]$ is an infinitely generated free Lie algebra with the generators
$[e_{j_{1}},[e_{j_{2}},[\cdots[e_{j_{s-1}},e_{j_{s}}]\cdots]],\qquad j_{1}\geqslant j_{2}\geqslant\cdots\geqslant j_{s-1}<j_{s},\,s\geqslant 2.$ (4.1)
@@
Denote by $\Delta$ the set of all pairs $(l,j)$ of positive integers such that $k\geqslant l>j\geqslant 1$. Here and throughout this section, when using the notation $l$ and $j$, it is always assumed that they form a pair $\delta\in\Delta$ and vice versa. It is natural to use polynomials in the operators $\operatorname{ad}e_{i}$. For $\delta\in\Delta$ and $\beta=(\beta_{1},\ldots,\beta_{l})\in\mathbb{Z}_{+}^{l}$ put
$g_{\delta,\beta}\!:=(\operatorname{ad}{e_{1}})^{\beta_{1}}\cdots(\operatorname{ad}{e_{l-1}})^{\beta_{l-1}}(\operatorname{ad}{e_{l}})^{\beta_{l}+1}(e_{j}).$ (4.2)
@@
where $T(V)$ is the tensor algebra of the vector space $V$. Note that (4.3) also holds for $k=1$ since in this case $\Delta$ is empty and so $V=0$.
@@
$\Psi_{\delta}\!:\mu_{1}^{\beta_{1}}\cdots\mu_{l}^{\beta_{l}}\mapsto(\operatorname{ad}e_{1})^{\beta_{1}}\cdots(\operatorname{ad}e_{l})^{\beta_{l}}$ (4.4)
@@
for $k\in\mathbb{N}$. (Here $[T](C^{\infty}(X))$ is the formal tensor algebra associated with the Fréchet space $C^{\infty}(X)$; see at the beginning of this section.)
@@
If an element $b$ of a real Banach algebra $B$ is of polynomial growth, then so is the operator $\operatorname{ad}b\!:B\to B$.
@@
where the hat means the Fourier transform. By Lemma 4.5, the operators $\operatorname{ad}b_{1},\ldots,\operatorname{ad}b_{k}$ are of polynomial growth. Therefore the function
$(s_{1},\ldots,s_{l})\mapsto\|\widehat{f}(s_{1},\ldots,s_{l})\,\exp\bigl{(}is_{1}\operatorname{ad}b_{1}\bigr{)}\cdots\exp\bigl{(}is_{l}\operatorname{ad}b_{l}\bigr{)}([b_{l},b_{j}])\|$
is absolutely integrable. Thus, the integral exists not only in the weak sense (of Gelfand and Pettis), but also in the strong sense (of Bochner) [23, p. 80, Theorem 3.7.4]. Applying Theorem 3.12 again, now to the tuple $(\operatorname{ad}b_{1},\ldots,\operatorname{ad}b_{k})$, we see that $\Psi_{\delta}$ is indeed an extension of the ordered calculus in (4.4). From the definition of $C^{\infty}(X)$ (see (4.5)), we have an extension of the map $V\to B$ to $C^{\infty}(X)\to B$.
@@
$\exp\bigl{(}is_{l}\operatorname{ad}b_{l}\bigr{)}([b_{l},b_{j}])=[\exp\bigl{(}is_{l}b_{l})\,b_{l}\exp\bigl{(}-is_{l}b_{l})\,,\,\exp\bigl{(}is_{l}b_{l})\,b_{j}\exp\bigl{(}-is_{l}b_{l})].$
By Theorem 2.4, each Banach algebra of polynomial growth is commutative modulo the radical. Therefore the integrand in (4.8) belongs to $\operatorname{Rad}B$ for all $s_{1},\ldots,s_{l}$. Since the radical of a Banach algebra is closed and the integral converges in the strong sense, we have that $\Psi_{\delta}(f)\in\operatorname{Rad}B$ for every $f\in C^{\infty}(X_{\delta})$. Therefore the image of $C^{\infty}(X)$ is also in $\operatorname{Rad}B$.
By Theorem 2.4, $\operatorname{Rad}B$ is nilpotent. Then we can apply the universal property for $[T](C^{\infty}(X))$ in Definition 4.1 and extend the continuous linear map $C^{\infty}(X)\to B$ constructed above to a continuous linear map from $[T](C^{\infty}(X))$ to $B$. The claim is proved.
@@
The main idea of the proof of Theorem 4.6 is the same as for Theorem 4.3 in [5] and Theorems 3.8 and 3.17 in this paper. We endow $U(\mathfrak{f}_{k})$ with the topology inherited from $C^{\infty}_{\mathfrak{f}_{k}}$ and construct a topologically injective homomorphism from $U(\mathfrak{f}_{k})$ to a product of algebras of the form $C^{\infty}(M,{\mathrm{T}}_{p})$, where $M$ is a manifold.
@@
$g_{\beta}=(\operatorname{ad}{e_{1}})^{\beta_{1}}(\operatorname{ad}{e_{2}})^{\beta_{2}+1}(e_{1}).$ (4.14)
@@
Then, applying inductively (4.13), (4.12) and (4.11) and using the fact that $\mathfrak{c}_{2}$ is a Lie ideal, we get the result. ∎